OpenStudy (katecc379):

Two competing internet companies have recently restructured their payment plans. Quick Net requires a $10 activation fee and charges $55 per month for their basic plan. Speedy Net requires no activation fee and charges $57 per month for their basic plan. When will the two plans be equal in cost?

OpenStudy (welshfella):

equation for Quick Net is P = 10 + 55n where P = total payment and n = number of months for Speedy net it is P = 57n

OpenStudy (welshfella):

when P is the same we have 10 + 55n = 57n solve for n

OpenStudy (welshfella):

do you know how to solve this ?

OpenStudy (katecc379):

no i dont

OpenStudy (katecc379):

can you explain how to so i understand

OpenStudy (welshfella):

OK you need to manipulate the equation so you isolate n first because LHS contains 55n we subtract 55n from it so it leaves the number 10. But we have to subtract 5n from the RHS also to keep the equation balanced. So we have 10 + 55n - 55n = 57n - 55n So as 55n - 55n = 0 we get:- 10 = 57n - 55n NOw 57n - 55n = 2n 10 = 2n now to get n from 2n we divide by 2 ( also we have to divide LHS by 2 10/2 = 2n/2 5 = n

OpenStudy (katecc379):

so after 5 months they would be equal?

OpenStudy (welshfella):

yes

OpenStudy (katecc379):

can you help with one more?

OpenStudy (katecc379):

@FaiqRaees

OpenStudy (faiqraees):

Can you draw the function g(x) on the graph?

OpenStudy (katecc379):

it wont let me its a quiz i just dont understand

OpenStudy (katecc379):

oh wait i get it i think haha

OpenStudy (faiqraees):

GOod

OpenStudy (katecc379):

would it be b ?

OpenStudy (faiqraees):

Yes

OpenStudy (katecc379):

@FaiqRaees

OpenStudy (katecc379):

sorry these are making no sense to me this lesson

OpenStudy (faiqraees):

They are asking you to find out the equation of the curves and then represent them in the form Equation of curve 1 = Equation of curve 2

OpenStudy (katecc379):

i dont understand

OpenStudy (katecc379):

would it be c?

OpenStudy (katecc379):

@FaiqRaees